OpenStudy (ams98):

Medal for explanation... Given the function f(x) = x3 + 2x2 − 3x − 5, what is resulting function when f(x) is shifted to the right 1 unit? f(x + 1) = x3 + 5x2 + 4x − 5 f(x) + 1 = x3 + 2x2 − 3x − 4 f(x) − 1 = x3 + 2x2 − 3x − 6 f(x − 1) = x3 − x2 − 4x − 1

OpenStudy (ams98):

f(x - 1) = (x - 1)^3 + 2(x - 1)^2 - 3(x - 1) - 5 = x^3 - 3x^2 + 3x - 1 + 2(x^2 - 2x + 1) - 3x + 3 - 5 = x^3 - x^2 - 4x - 1 ??

OpenStudy (perl):

want me to expand it?

OpenStudy (perl):

x^3-x^2-4*x-1

OpenStudy (perl):

yes f(x-1) is the correct answer

OpenStudy (squirrels):

ashley pls